# Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior

### v4

@@ VIII Discussion and Conclusion
Under the concave game condition, from Theorems 5 and 9, if $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$, we have
$x^{*}$ $=-r_{\mathrm{x}},y^{*}=-r_{\mathrm{y}},x_{\mathrm{cen}}^{*}=-r_{\mathrm{x}},y_{\mathrm{cen}}^{*}=-r_{\mathrm{y}},$ (59)
and change the critical point from $-r_{\mathrm{x}}$ to $r_{\mathrm{x}}$ will get the solution when $S_{\mathrm{b},1}<S_{\mathrm{b},2}$. Thus, $x^{*}=x_{\mathrm{cen}}^{*},y^{*}=y_{\mathrm{cen}}^{*}$ and $P=1$, indicating concave game condition is equivalent to centralized condition.
From Theorem 10, we know the PoA under quasiconcave and non-concave game conditions can be written as (57b), and nominator minus denominator is $(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}\geq 0$. Thus, $P\geq 1$ under these two conditions, indicating quasiconcave and non-concave games always cause higher (or equal) anarchy than concave games.
Then Given fixed $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$, the PoA (57b) is only affected by the solution structure $x^{*},y^{*}$. Note that $X_{1},X_{2},Y_{1},Y_{2}$ could be variant such that $X_{1}+X_{2}+Y_{1}+Y_{2}=2S$, which may cause quasiconcave or non-concave game condition, i.e., all agents satisfy the following or not,
@@
We basically fixed the other parameters that appeared in the PoA expression of (57b) to explicitly show the influence of the game type change. According to Theorem 5, we know both quasiconcave and non-concave games balance the system demand in the two period periods, i.e., $x^{*}+y^{*}=b$. For the quasiconcave game, $x^{*}=b_{\mathrm{x}},y^{*}=b_{\mathrm{y}}$, and satisfy the (60), while for non-concave game, $x^{*}=\pm r_{\mathrm{x}},y^{*}=b\mp r_{\mathrm{x}}$, and $b_{\mathrm{x}}>r_{\mathrm{x}}$ or $b_{\mathrm{x}}<-r_{\mathrm{x}}$. Suppose $x^{*}=r_{\mathrm{x}},y^{*}=b-r_{\mathrm{x}}$, then the condition is $b_{\mathrm{x}}>r_{\mathrm{x}}$ and $b_{\mathrm{y}}<b-r_{\mathrm{x}}$, indicating
@@
If $x^{*}=-r_{\mathrm{x}},y^{*}=b+r_{\mathrm{x}}$, the situation is similar and we omit the redundant math here. Thus, the quasiconcave game causes higher PoA than the non-concave game under fixed $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$. These finish the proof of the Theorem. ∎
@@
Overview of the proof: We derive two lemmas to show the existence and uniqueness of NE in the multi-agent CP game $G^{\prime}$, respectively. We first show the CP game $G^{\prime}$ exists NE (Lemma 17) and show the NE exists in Lemma 17 is unique (Lemma 18). We first introduce two concepts.
(1) Payoff security (Reny [Reny, 1999]). Agent $i$ can secure the payoff $f_{i}(x_{i},x_{-i})-\epsilon\in\mathbb{R}$ at $x_{i},x_{-i}\in\mathcal{X}$ iff for every $\epsilon>0$, there exist a $\hat{x}_{i}\in\mathcal{X}_{i}$ such that $f_{i}(\hat{x}_{i},x_{-i}^{\prime})\geq f_{i}(x_{i},x_{-i})-\epsilon$ for every $x_{-i}^{\prime}$ in some neighborhood of $x_{-i}$. Furthermore, we say that a game $G^{\prime}$ is payoff secure iff every agent $i$ can secure payoff for every $x_{i}\in\mathcal{X}_{i}$.
(2) Diagonally strictly concave (Rosen [Rosen, 1965]). Define the pseudo-gradient of the sum of all agents’ payoff functions $\sum_{i\in N}f_{i}(x_{i},x_{-i})$ and total differential operator $\nabla$ as
@@
Among them, payoff security means every agent can secure a payoff value in any strategy profile if they have a strategy that provides at least this value, even if other agents slightly change their strategies. We also have the sufficient conditions for diagonally strictly concave function from Rosen [Rosen, 1965], namely that the symmetric matrix $G(x_{i},x_{-i})+G^{T}(x_{i},x_{-i})$ be negative definite for $x_{i},x_{-i}\in\mathcal{X}$, where $G(x_{i},x_{-i})$ is the Jacobian of $F(x_{i},x_{-i})$ with respect to all $x_{i}$. We then introduce the existence lemma.
@@
Existence. The multi-agent two-period quasiconcave CP game $G^{\prime}$ as described in (19) and (20b), has a pure-strategy NE $(x_{i}^{*},x_{-i}^{*})$ as defined in (21).
@@
According to Reny [Reny, 1999], a compact, convex, bounded, and quasiconcave game has a pure strategy NE if the sum of the agent’s payoff functions is u.s.c. as defined in Definition 1 in the whole strategy set and the game is payoff security as defined above. Thus, we focus on (a) the sum of the agent’s payoff functions, i.e., $\sum_{i\in N}f_{i}(x_{i},x_{-i})$, is u.s.c. in $x_{i},x_{-i}\in\mathcal{X}$ and (b) the game $G^{\prime}$ is payoff secure, to prove the existence of pure-strategy NE.
@@
$\sum_{i\in N}f_{i}(x_{i},x_{-i})$ $=-\pi\sum_{i\in N}(X_{i,1}+x_{i})I(S_{1}-S_{2})-\pi\sum_{i\in N}(X_{i,2}-x_{i})I(S_{2}-S_{1})-\sum_{i\in N}\alpha_{i}x_{i}^{2}$
$=\pi S_{1}I(S_{1}-S_{2})-\pi S_{2}I(S_{2}-S_{1})-\sum_{i\in N}\alpha_{i}x_{i}^{2}.$ (63)
This function is u.s.c. according to Definition 1 due to $\pi S_{1}=\pi S_{2}$ when $S_{1}=S_{2}$.
(b). We then show that each agent is payoff secure by analyzing their strategies. Given any strategy pair $(x_{i},x_{-i})$ and for any $\epsilon>0$, agents $i,-i$ can secure the payoff $f_{i}(x_{i},x_{-i})-\epsilon$ and $f_{-i}(x_{i},x_{-i})-\epsilon$. Note that here $-i$ denotes all agents but $i$.
Suppose the CP period is 1, i.e., $S_{1}\geq S_{2}$, if one of the agents among $-i$ slightly increases $x_{-i}$, the CP period is still 1, and agent $i$’s payoff is the same if agent $i$ keeps the strategy profile $x_{i}$; if one of the agents among $-i$ slightly decrease $x_{-i}$, CP period is possible to change, but agent $i$ can secure the payoff $f_{i}(x_{i},x_{-i})-\epsilon$ by slightly reducing its strategy $x_{i}$. Basically, there exist a $\delta>0$ small enough that satisfy $f_{i}(x_{i}+\delta,x_{-i}^{\prime})>f_{i}(x_{i},x_{-i})-\epsilon$ for all $S_{1}^{\prime}=X_{i,1}+x_{i}+\delta+\sum_{-i\in N}(X_{-i,1}+x_{-i}^{\prime})>S_{2}$. Thus, agents $i$ is payoff secure; a similar analysis can be used for all agents $-i$ by altering $i$, and CP period 2 also follows the same analysis. Thus, the CP game is payoff secure.
According to the Reny [Reny, 1999], the CP game has a pure-strategy NE. ∎
@@
According to Rosen’s method [Rosen, 1965], if the sum of the agent’s payoff function is diagonally strictly concave as defined in Definition 1, then the equilibrium point described in Lemma 17 is unique if the constraints in the strategy set are concave, i.e., $\mathcal{X}=\{x|h(x)\geq 0\}$, where $h(x)$ is a concave function.
@@
The Hessian $H$ with respect to $x_{i}$ is $G=G^{T}=-2\text{diag}(\alpha_{1},\cdots,\alpha_{N})$, where $\text{diag}(\cdot):\mathbb{R}^{N}\to\mathbb{R}^{N\cdot N}$. This shows that $G+G^{T}$ is obviously negative definite, thus, we know the sum of the agent’s payoff function is diagonally strictly concave. This means, according to Rosen’s method, there is a unique NE in the strategy set $\mathcal{X}_{\mathrm{cp}1}$. Following the same analysis, we can conclude there is a unique NE in the strategy set $\mathcal{X}_{\mathrm{cp}2}$.
Now, the difficult part of proving this Lemma lies in the switching between two strategy sets. Thus, we separate three cases according to the CP period to analyze the NE.
  • Case (i): CP period is always 1, i.e., $S_{1}(x_{i},x_{-i})\geq S_{2}(x_{i},x_{-i}),\forall x_{i},x_{-i}\in\mathcal{X}$;
  • Case (ii): CP period is always 2, i.e., $S_{1}(x_{i},x_{-i})<S_{2}(x_{i},x_{-i}),\forall x_{i},x_{-i}\in\mathcal{X}$.
  • Case (iii): CP period changes during the gaming process, i.e., $\exists x_{i}^{\prime},x_{-i}^{\prime},x_{i}^{\prime\prime},x_{-i}^{\prime\prime}\in\mathcal{X},S_{1}(x_{i}^{\prime},x_{-i}^{\prime})-S_{2}(x_{i}^{\prime},x_{-i}^{\prime})\geq 0,S_{1}(x_{i}^{\prime\prime},x_{-i}^{\prime\prime})-S_{2}(x_{i}^{\prime\prime},x_{-i}^{\prime\prime})<0$;
We first show Cases (i) and (ii) can’t exist under the quasiconcave conditions (20b) because only one period could be CP period during the entire solution process and the game is concave. Taking case (i) for example, mathematically, we have, $S_{1}(x_{i},x_{-i})\geq S_{2}(x_{i},x_{-i}),\forall x_{i},x_{-i}\in\mathcal{X}=\mathcal{X}_{\mathrm{cp}1}$, i.e.,
$S_{1}(x_{i},x_{-i})$ $=\sum_{i\in N}(X_{i,1}+x_{i})\geq\sum_{i\in N}(X_{i,2}-x_{i})=S_{2}(x_{i},x_{-i}),$ (65a)
$\frac{\sum_{i\in N}(X_{i,2}-X_{i,1})}{2}=b\leq\sum_{i\in N}x_{i},$ (65b)
which aligns with the conditions of the concave game. The same analysis can be applied to case (ii), where the CP period is always 2. Thus, we know both cases (i) and (ii) can’t exist under the quasiconcave conditions.
We then analyze Case (iii) and start with the two-agent setting with agent x and y. We first note that the CP charge, either $\pi(X_{1}+x)$ or $\pi(X_{2}-x)$, is relatively greater compared with penalty $\alpha_{\mathrm{x}}x^{2}$ for all agents because the CP period can change in the case (iii). From the best response perspective, suppose the CP period is 1, i.e., $S_{1}=S_{2}+\delta$, where $\delta$ is an infinitesimal number. Now, if agent x benefits by reducing $x$, which changes the CP period to 2. This indicates the demand of agent x in period 1 is greater than period 2, i.e., $X_{1}+x\geq X_{2}-x$. Now if agent y’s demand in period 1 is less than period 2, i.e., $Y_{1}+y<Y_{2}-y$, then CP period change from 1 to 2 harms agent y’s benefit, so agent y wants to push the CP period back to 1 by increasing $y$; if $Y_{1}+y\geq Y_{2}-y$, CP period changes to 2 means agent y can benefit by shifting $y$ away from period 2, i.e., still increase $y$. This means the best response of agent y is always adversarial to that of agent x.
This shows agents are fully competitive; basically, if agent x can benefit from changing strategy, that strategy harms another agent’s benefit. This is similar to the zero-sum game setting. Thus, we can apply the following min-max formulation to analyze the best responses,
where $y^{*}$ is the optimal solution of agent y following the same structure, and the first and second case corresponds to CP periods 1 and 2. By applying the first-order optimality condition, the optimal solution of $x$ is
The optimal solution for agent y follows the same structure
According to (5), $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}$, $-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}}$, we know if $S_{1}\geq S_{2}$, when $x^{*},y^{*}$ reach optimal solution, $y^{*}=\max\{b-x^{*},b_{\mathrm{y}}\}$ and $x^{*}=\max\{b-y^{*},b_{\mathrm{x}}\}$. Then, according to (66b) and (66c), the only solutions are $x^{*}=b_{\mathrm{x}},\ y^{*}=b_{\mathrm{y}}$, and the conditions are
$X_{2}+Y_{2}-x^{*}-y^{*}=S_{2}=X_{1}+Y_{1}+x^{*}+y^{*}=S_{1}.$ (67)
@@
Then, we use the two-agent solution to analyze the solution in the $N$-agent setting. Given that all agents are capable, define a partition of set $N$ into two disjoint subsets, denoted as $N_{\mathrm{a}},N_{\mathrm{b}}$, such that $N_{\mathrm{a}}\cup N_{\mathrm{b}}=N,N_{\mathrm{a}}\cap N_{\mathrm{b}}=\emptyset$, and the demand for these two subsets is unequal in at least one period, i.e.,
$\{\sum_{i\in N_{\mathrm{a}}}(X_{i,1}+x_{i})\neq\sum_{i\in N_{\mathrm{b}}}(X_{i,1}+x_{i})\}\cup\{\sum_{i\in N_{\mathrm{a}}}(X_{i,2}-x_{i})\neq\sum_{i\in N_{\mathrm{b}}}(X_{i,2}-x_{i})\}.$ (68)
@@
Then treat these two sets as two aggregated agents with strategy $x_{\mathrm{a}},x_{\mathrm{b}}$ and apply the two-agent solutions, we know the two sets will balance their demand over the two periods, i.e., the following is true for sets $N_{\mathrm{a}},N_{\mathrm{b}}$,
$x_{\mathrm{a}}^{*}=\frac{\sum_{i\in N_{\mathrm{a}}}X_{i,2}-\sum_{i\in N_{\mathrm{a}}}X_{i,1}}{2}=\frac{\sum_{i\in N_{\mathrm{a}}}(X_{i,2}-X_{i,1})}{2}=\sum_{i\in N_{\mathrm{a}}}b_{i}=\sum_{i\in N_{\mathrm{a}}}x_{i}^{*}.$ (69)
Then, applying the same partition to $N_{\mathrm{a}},N_{\mathrm{b}}$ to get the subsets $N_{\mathrm{a,a}},N_{\mathrm{a,b}}$ and $N_{\mathrm{b,a}},N_{\mathrm{b,b}}$, and applying the two-agent solutions, we still know the two subsets will balance demand over the two periods, i.e., $\sum_{i\in N_{\mathrm{a,a}}}x_{i}^{*}=\sum_{i\in N_{\mathrm{a,a}}}b_{i}$ for sets $N_{\mathrm{a,a}},N_{\mathrm{a,b}},N_{\mathrm{b,a}},N_{\mathrm{b,b}}$. Iterating applying the partition to each subset to get corresponding two subsets, the subset will finally become a singleton and we have $x_{i}^{*}=b_{i},i\in N$, and
$\sum_{i\in N}(X_{i,2}-x_{i}^{*})=S_{2}=\sum_{i\in N}(X_{i,1}+x_{i}^{*})=S_{1}.$ (70)
@@
This shows the NE will always be obtained when $S_{1}=S_{2}$ under the quasiconcave condition, i.e., all agents’ are capable as described in Proposition 4. This also means during the game, periods 1 and 2 are active interactively, and the agent’s strategy set switches between $\mathcal{X}_{\mathrm{cp}1}$ and $\mathcal{X}_{\mathrm{cp}2}$ and finally converges at the connecting point of both strategy set $S_{1}=S_{2}$. This means there is a unique equilibrium point in the entire strategy set $\mathcal{X}$. Indeed, all agents will minimize the payment associated with $S_{1},S_{2}$ imbalance as any imbalance results in a significant CP charge change caused by the opponent’s strategy.
@@
We first show the quasiconcave multi-agent CP game $G^{\prime}$ as described in (19) and (20b) has pure-strategy NE based on Lemma 17. Then, we prove the NE existed in Lemma 17 is unique based on Lemma 18 and the unique NE is obtained as described in (23). ∎
@@
Given $S_{\mathrm{b},1}<S_{\mathrm{b},2}$. We first assume two virtual agents as CP agent and non-CP agent. We denote the strategy as $x_{\mathrm{cp}}$, baseline demand as $X_{\mathrm{cp},2}=\sum_{i\in N_{\mathrm{cp}}}X_{i,2},X_{\mathrm{cp},1}=\sum_{i\in N_{\mathrm{cp}}}X_{i,1}$, balance point $b_{\mathrm{cp}}=(X_{\mathrm{cp},2}-X_{\mathrm{cp},1})/2$, and critical point $r_{\mathrm{cp}}=\sum_{i\in N_{\mathrm{cp}}}r_{i}$ for CP agent, and change the subscript to $\mathrm{ncp}$ for non-CP period agent. Note that for CP agent $X_{\mathrm{cp},1}<X_{\mathrm{cp},2}$ (the same definition with CP-period agent) and non-CP agent vice verse, we also have $b_{\mathrm{cp}}+b_{\mathrm{ncp}}=b$.
Then, from Theorem 5, we know these two virtual agents will balance the system demand at equilibrium. Depending on which agent reaches the critical point first, the best strategy for one agent is the critical point and the other is to balance the left unbalanced system demand. Suppose the CP agent reaches the critical point first, i.e., the demand shifting of the non-CP agent cannot be offset by the CP agent, meaning that the CP agent can’t flatten the system demand given the non-CP agent’s shifting even if the CP agent shifts the maximum amount of demand. We can express the condition as
$r_{\mathrm{cp}}<b_{\mathrm{cp}},-r_{\mathrm{ncp}}-r_{\mathrm{cp}}\leq b\leq r_{\mathrm{ncp}}+r_{\mathrm{cp}},$ (71a)
and the NE between these two virtual agents is
@@
We then aggregate CP-period agents and non-CP-period agents into the CP-period agent set and non-CP-period agent set and fix the agent in each agent set, as the CP-period agent and non-CP-period agent are defined based on their baseline demand conditions. Obviously, the baseline demand, balance point, and critical point are equivalent between the agents set and the virtual agent. However, they are not totally equivalent because their strategies are different. The strategy for the CP-period agent set is determined by each agent $x_{i}^{*},i\in N_{\mathrm{cp}}$, and the non-CP-period agent set follows the same idea. Note that the same baseline demand and critical point indicate their best response rationale is the same; basically, the virtual CP agent benefits by shifting demand away from the CP period, and we have the same for the agents in the CP-period agent set.
From Theorem 5, we know each agent’s maximum shifting capacity is $b_{i}$, i.e., $x_{i}^{*}=\min\{r_{i},b_{i}\}$, and from the condition (71a), we know although the CP-period agent set and virtual CP agent have the same conditions, CP-period agent set can’t reach the best strategy of the virtual CP agent due to the following

### v5

@@ VIII Discussion and Conclusion
Under the concave game condition, from Theorems 1 and 4, we have
$x^{*}$ $=r_{\mathrm{x}},y^{*}=r_{\mathrm{y}},x_{\mathrm{cen}}^{*}=r_{\mathrm{x}},y_{\mathrm{cen}}^{*}=r_{\mathrm{y}},$ (60)
which shows $x^{*}=x_{\mathrm{cen}}^{*},y^{*}=y_{\mathrm{cen}}^{*}$ and $P=1$, meaning that the concave game condition is equivalent to centralized model.
From Theorem 5, we know the efficiency loss under quasiconcave and non-concave game conditions can be written as (58b), and nominator minus denominator is $(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}\geq 0$. Thus, $P\geq 1$ under these two conditions, indicating quasiconcave and non-concave games always cause higher (or equal) anarchy than concave games.
Then Given fixed $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$, the efficiency loss (58b) is only affected by the solution structure $x^{*},y^{*}$. Note that $X_{1},X_{2},Y_{1},Y_{2}$ could be variant such that $X_{1}+X_{2}+Y_{1}+Y_{2}=2S$, which may cause quasiconcave or non-concave game condition, i.e., whether both agents satisfy the following conditions
@@
We basically fixed the other parameters that appeared in the efficiency loss expression of (58b) to explicitly show the influence of the game type change. According to Theorem 1, we know both quasiconcave and non-concave games balance the system demand over the two periods, i.e., $x^{*}+y^{*}=b$. For the quasiconcave game, $x^{*}=b_{\mathrm{x}},y^{*}=b_{\mathrm{y}}$, and satisfy the (61), while for non-concave game, $x^{*}=r_{\mathrm{x}},y^{*}=b-r_{\mathrm{x}}$, and $b_{\mathrm{x}}>r_{\mathrm{x}}$. Suppose $x^{*}=r_{\mathrm{x}},y^{*}=b-r_{\mathrm{x}}$, then the condition is $b_{\mathrm{x}}>r_{\mathrm{x}}$ and $b_{\mathrm{y}}<b-r_{\mathrm{x}}$, indicating
@@
If $x^{*}=b\pm r_{\mathrm{y}},y^{*}=\mp r_{\mathrm{y}}$, the situation is similar and we omit the redundant math here. Thus, the quasiconcave game causes higher efficiency loss than the non-concave game under fixed $\pi,S,\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}>0$. These finish the proof of the Theorem. ∎
@@
Overview of the proof: We derive two lemmas to show the existence and uniqueness of NE in the multi-agent CP shaving game $G^{\prime}$, respectively. We first show the CP shaving game $G^{\prime}$ exists NE (Lemma 3) and show the NE exists in Lemma 3 is unique (Lemma 4). We first introduce two concepts.
(1) Payoff security (Reny [Reny,, 1999]). Agent $i$ can secure the payoff $f_{i}(x_{i},x_{-i})-\epsilon\in\mathbb{R}$ at $x_{i},x_{-i}\in\mathcal{X}$ iff for every $\epsilon>0$, there exist a $\hat{x}_{i}\in\mathcal{X}_{i}$ such that $f_{i}(\hat{x}_{i},x_{-i}^{\prime})\geq f_{i}(x_{i},x_{-i})-\epsilon$ for every $x_{-i}^{\prime}$ in some neighborhood of $x_{-i}$. Furthermore, we say that a game $G^{\prime}$ is payoff secure iff every agent $i$ can secure payoff for every $x_{i}\in\mathcal{X}_{i}$.
(2) Diagonally strictly concave (Rosen [Rosen,, 1965]). Define the pseudo-gradient of the sum of all agents’ payoff functions $\sum_{i\in N}f_{i}(x_{i},x_{-i})$ and total differential operator $\nabla$ as
@@
Among them, payoff security means every agents can secure a payoff value in any strategy profile if they have a strategy that provides at least this value, even if other agents slightly change their strategies. We also have the sufficient conditions for a diagonally strictly concave function from Rosen [Rosen,, 1965], namely that the symmetric matrix $G(x_{i},x_{-i})+G^{T}(x_{i},x_{-i})$ be negative definite for $x_{i},x_{-i}\in\mathcal{X}$, where $G(x_{i},x_{-i})$ is the Jacobian of $F(x_{i},x_{-i})$ with respect to all $x_{i}$. We then introduce the existence lemma.
@@
Existence. The multi-agent two-period quasiconcave CP shaving game $G^{\prime}$ as described in (19) and (20b), has a pure-strategy NE $(x_{i}^{*},x_{-i}^{*})$ as defined in (21).
@@
According to Reny [Reny,, 1999], a compact, convex, bounded, and quasiconcave game has a pure strategy NE if the sum of the agent’s payoff functions is u.s.c. as defined in Definition 1 across the entire strategy set and if the game satisfies payoff security as defined above. Intuitively, the game is compact, convex, and bounded, meaning that the strategy spaces $\mathcal{X}i,i\in N$ are compact and convex, while the payoff functions $f{i},i\in N$ remain bounded due to the quadratic shifting penalty term. Thus, to establish the existence of a pure-strategy NE, we focus on verifying two key conditions: (a) the sum of the agents’ payoff functions, $\sum_{i\in N}f_{i}(x_{i},x_{-i})$, is u.s.c. for all $x_{i},x_{-i}\in\mathcal{X}$; and (b) the game $G^{\prime}$ satisfies the payoff security condition.
@@
$\sum_{i\in N}f_{i}(x_{i},x_{-i})$ $=-\pi\sum_{i\in N}(X_{i,1}+x_{i})I(x_{i},x_{-i})-\pi\sum_{i\in N}(X_{i,2}-x_{i})(1-I(x_{i},x_{-i}))-\sum_{i\in N}\alpha_{i}x_{i}^{2}$
$=-\pi S_{1}I(x_{i},x_{-i})-\pi S_{2}(1-I(x_{i},x_{-i}))-\sum_{i\in N}\alpha_{i}x_{i}^{2}.$ (64)
This function is u.s.c. as per Definition 1 because, on the switching surface where $S_{1}=S_{2}$, the cost term satisfies $\pi S_{1}=\pi S_{2}$.
(b). We now demonstrate that each agent is payoff secure by analyzing their strategic choices. Given any strategy pair $(x_{i},x_{-i})$ and for any $\epsilon>0$, agents $i,-i$ can secure payoffs of at least $f_{i}(x_{i},x_{-i})-\epsilon$ and $f_{-i}(x_{i},x_{-i})-\epsilon$, respectively. Note that here $-i$ denotes all agents but $i$.
Suppose the CP period is 1, meaning $S_{1}\geq S_{2}$. If any agent among $-i$ slightly increases their demand shift $x_{-i}$, the CP period remains 1, and agent $i$’s payoff remains unchanged if they maintain their strategy $x_{i}$. Conversely, if any agent among $-i$ slightly reduces $x_{-i}$, the CP period may switch to 2. However, agent $i$ can still secure a payoff of at least $f_{i}(x_{i},x_{-i})-\epsilon$ by making a slight reduction in their strategy $x_{i}$. Specifically, there exists a sufficiently small $\delta>0$ such that $f_{i}(x_{i}+\delta,x_{-i}^{\prime})>f_{i}(x_{i},x_{-i})-\epsilon$ for all cases where the updated demand condition $S_{1}^{\prime}=X_{i,1}+x_{i}+\delta+\sum_{-i\in N}(X_{-i,1}+x_{-i}^{\prime})>S_{2}$ still holds. This confirms that agent $i$ is payoff secure. A similar argument applies symmetrically to all other agents $-i$ and the same reasoning extends to the case where the CP period is 2. Thus, the CP shaving game is payoff secure.
By Reny’s theorem [Reny,, 1999], we conclude that the CP shaving game has a pure-strategy NE. ∎
@@
According to Rosen’s method [Rosen,, 1965], if the sum of the agents’ payoff functions is diagonally strictly concave, as defined earlier, then the equilibrium point established in Lemma 3 is unique, provided that the constraints in the strategy set are concave. That is, if the strategy space is defined as $\mathcal{X}=\{x|h(x)\geq 0\}$, where $h(x)$ is a concave function, the uniqueness of the equilibrium follows.
@@
The Hessian $H$ with respect to $x_{i}$ is $G=G^{T}=-2\text{diag}(\alpha_{1},\cdots,\alpha_{N})$, where $\text{diag}(\cdot):\mathbb{R}^{N}\to\mathbb{R}^{N\cdot N}$. This shows that $G+G^{T}$ is clearly negative definite, confirming that the sum of the agents’ payoff functions is diagonally strictly concave. By Rosen’s method, this ensures the existence of a unique NE within the strategy set $\mathcal{X}_{\mathrm{cp}1}$. Applying the same reasoning, we can similarly conclude that a unique NE exists within the strategy set $\mathcal{X}_{\mathrm{cp}2}$.
The primary challenge in proving this lemma lies in analyzing the transition between the two strategy sets. Specifically, we need to show the existence of strategies $x_{i}^{\prime},x_{-i}^{\prime},x_{i}^{\prime\prime},x_{-i}^{\prime\prime}\in\mathcal{X}$ such that $S_{1}(x_{i}^{\prime},x_{-i}^{\prime})-S_{2}(x_{i}^{\prime},x_{-i}^{\prime})\geq 0,S_{1}(x_{i}^{\prime\prime},x_{-i}^{\prime\prime})-S_{2}(x_{i}^{\prime\prime},x_{-i}^{\prime\prime})<0$.
@@
Then, we leverage the two-agent solution from Lemma 1 to extend our analysis to the $N$-agent setting. Given that all agents are capable, we define a partition of the set $N$ into two disjoint subsets, denoted as $N_{\mathrm{a}},N_{\mathrm{b}}$, such that $N_{\mathrm{a}}\cup N_{\mathrm{b}}=N,N_{\mathrm{a}}\cap N_{\mathrm{b}}=\emptyset$. This partitioning ensures that the total demand of these two subsets is unequal in at least one period, i.e.,
$\{\sum_{i\in N_{\mathrm{a}}}(X_{i,1}+x_{i})\neq\sum_{i\in N_{\mathrm{b}}}(X_{i,1}+x_{i})\}\cup\{\sum_{i\in N_{\mathrm{a}}}(X_{i,2}-x_{i})\neq\sum_{i\in N_{\mathrm{b}}}(X_{i,2}-x_{i})\}.$ (66a)
@@
Then, we treat these two sets as two aggregated agents with strategies $x_{\mathrm{a}},x_{\mathrm{b}}$ and apply the two-agent solutions from Lemma 1 again. By doing so, we establish that these two sets will balance their total demand over the two periods, i.e., the following is true for sets $N_{\mathrm{a}},N_{\mathrm{b}}$,
$x_{\mathrm{a}}^{*}=\frac{\sum_{i\in N_{\mathrm{a}}}X_{i,2}-\sum_{i\in N_{\mathrm{a}}}X_{i,1}}{2}=\frac{\sum_{i\in N_{\mathrm{a}}}(X_{i,2}-X_{i,1})}{2}=\sum_{i\in N_{\mathrm{a}}}b_{i}=\sum_{i\in N_{\mathrm{a}}}x_{i}^{*}.$ (66b)
Then, applying the same partitioning process to $N_{\mathrm{a}},N_{\mathrm{b}}$, we obtain the subsets $N_{\mathrm{a,a}},N_{\mathrm{a,b}}$ and $N_{\mathrm{b,a}},N_{\mathrm{b,b}}$. Applying the two-agent solutions again, we establish that these subsets will also balance their demand over the two periods, i.e., $\sum_{i\in N_{\mathrm{a,a}}}x_{i}^{*}=\sum_{i\in N_{\mathrm{a,a}}}b_{i}$ for each subset. Iteratively applying this partitioning process to each subset, we continue refining them into smaller disjoint sets. Eventually, as the subsets reduce to singletons, we obtain the final equilibrium solution for each individual agent: $x_{i}^{*}=b_{i},i\in N$ and
$\sum_{i\in N}(X_{i,2}-x_{i}^{*})=S_{2}=\sum_{i\in N}(X_{i,1}+x_{i}^{*})=S_{1}.$ (66c)
@@
This means during the game, periods 1 and 2 are active interactively, and the agent’s strategy set switches between $\mathcal{X}_{\mathrm{cp}1}$ and $\mathcal{X}_{\mathrm{cp}2}$ and finally converges at the connecting point of both periods, i.e., $S_{1}=S_{2}$. This means there is a unique equilibrium point in the entire strategy set $\mathcal{X}$.
@@
We first show the quasiconcave multi-agent CP shaving game $G^{\prime}$ as described in (19) and (20b) has pure-strategy NE based on Lemma 3. Then, we prove the NE existed in Lemma 3 is unique based on Lemma 4 and the unique NE is obtained as described in (23). ∎
@@
We first assume two virtual agents as CP agent and non-CP agent. We denote their strategy as $x_{\mathrm{cp}}$, baseline demand as $X_{\mathrm{cp},2}=\sum_{i\in N_{\mathrm{cp}}}X_{i,2},X_{\mathrm{cp},1}=\sum_{i\in N_{\mathrm{cp}}}X_{i,1}$, balance point as $b_{\mathrm{cp}}=(X_{\mathrm{cp},2}-X_{\mathrm{cp},1})/2$, and critical point as $r_{\mathrm{cp}}=\sum_{i\in N_{\mathrm{cp}}}r_{i}$ for CP agent, and change the subscript to $\mathrm{ncp}$ for non-CP period agent. Note that for CP agent $X_{\mathrm{cp},1}<X_{\mathrm{cp},2}$ (the same definition with individual CP-period agent) and non-CP agent vice verse, we also have $b_{\mathrm{cp}}+b_{\mathrm{ncp}}=b$.
Then, from Theorem 1, we know that these two virtual agents will balance the system demand at equilibrium. The equilibrium outcome depends on which agent reaches its critical point first. If the CP agent reaches its critical point first, it indicates that the non-CP agent’s demand shifting cannot be fully offset by the CP agent. In other words, even if the CP agent shifts the maximum possible amount of demand, it is still unable to completely flatten the system demand across the two periods. In this case, the optimal strategy is for the CP agent to shift demand up to its critical point, while the non-CP agent balances the remaining unbalanced system demand. We can express the condition as
$r_{\mathrm{cp}}<b_{\mathrm{cp}},0\leq b\leq r_{\mathrm{ncp}}+r_{\mathrm{cp}},$ (67a)
and the NE of these two virtual agents are
@@
We then aggregate the CP-period agents and non-CP-period agents into the CP-period agent set and non-CP-period agent set, respectively, while keeping the classification of each agent fixed based on their baseline demand conditions. The baseline demand, balance point, and critical point of each agent set are equivalent to those of the corresponding virtual agent. However, they are not entirely identical, as their strategy structures differ. The strategy for the CP-period agent set is determined by the individual decisions of its agents, denoted as $x_{i}^{*},i\in N_{\mathrm{cp}}$, and similarly for the non-CP-period agent set. Despite this distinction, having the same baseline demand and critical point ensures that their best response rationales are aligned. Essentially, just as the virtual CP agent benefits from shifting demand away from the CP period, the agents within the CP-period agent set follow the same strategic behavior.
By the model formulation, each agent’s maximum shifting capacity is $b_{i}$, leading to the best strategy $x_{i}^{*}=\min{r_{i},b_{i}},i\in N_{\mathrm{cp}}$. Furthermore, from condition (67a), we observe that although the CP-period agent set and the virtual CP agent share the same baseline demand conditions, the CP-period agent set cannot achieve the optimal strategy of the virtual CP agent due to the following